# Many-Body Localization in the Age of Classical Computing

### v2

@@ III.2 Ergodicity breaking: single particle systems and beyond
While ETH has been observed numerically in a wide class of many-body systems, there is no proof of it for a generic, local Hamiltonian, and provable counterexamples of fine-tuned, integrable, Hamiltonians can be constructed. In this sense, the situation is similar to classical mechanics’ Ergodic Hypothesis, which is widely believed to be true for generic Hamiltonian dynamics (for a sufficiently large number of degrees of freedom), but could only be proved for some realistic systems [29, 607].
Anderson localization [27] is often referred to as a counter-example to ETH, which has no counterpart in classical mechanics. However, this interpretation requires some care, since the Anderson model, invoked to model Anderson localization on a lattice, is a quadratic model, [^2] Quadratic Hamiltonians are sums of bilinear forms of creation and annihilation operators, see Eq. (5) for one example. and it can be proven that many-body eigenstates of quadratic models can not exhibit ETH in the many-particle Hilbert space [729]. However, recent work has formulated the notion of single-particle eigenstate thermalization [730], which can be seen as the analogue of (3), defined for the matrix elements in single-particle eigenstates of quadratic models, and a natural consequence of the Berry conjecture [62]. Within this framework, Anderson insulator should be seen as an example of the breakdown of single-particle eigenstate thermalization.
@@
where $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) are fermionic creation (anihilation) operators creating particle at site $i$, the sum over neighboring lattice sitesis denoted by $\braket{i,j}$, and $\epsilon_{i}$ are on-site energies. In case the on-site energies $\epsilon_{i}$ are a random uncorrelated numbers, the model in Eq. (5) is referred to as the Anderson model. If the variance of $\epsilon_{i}$ is large enough (with respect to the hopping rate), it turns out that each particle stays confined close to its original position, and the system avoids the process of reaching thermal equilibrium. It goes without saying that there is no conduction in the system and the system is an insulator. In this case, the single-particle version of the ETH ansatz (3) does not apply [730], and the long-time value of a local observable such as site occupation is not recovered by the microcanonical average (4). This gives rise to the absence of transport and several mathematical proofs of localization at large disorder exist [234, 213, 12, 85] [^3] It is curious however that no proof of absence of localization at small disorder exist for the Anderson model except for the Bethe lattice [13]..
In spite of the above arguments, it could not be ruled out that interactions would destroy the one-electron picture of Anderson localization and a small conductivity will be born, possibly even exponentially small in the inverse of the interaction strength (measured in units of the local density of states). In fact, a mechanism in which a small interaction between electrons and phonons can induce a non-zero conductivity via the phonon-assisted hopping was already pointed out by Mott [432]. However, crucially, the phonons considered by Mott are supposed to be a good, ergodic, bath not perturbed by the presence of localized electrons. But can a bath of localized electrons provide the necessary frequencies for a conduction probe electron to move for macroscopic distances? In other words, could electrons act like Mott’s phonon bath? Fleishman and Anderson [208] already in the 1980’s proposed, on the basis of perturbation theory (and some general intuition), that, at small enough temperatures and weak interactions, localization would be preserved.
@@ III.3 MBL phase: a definition
Twenty six years after Fleishman and Anderson, [51, 244] presented a thorough analysis of the perturbation theory, where they found that at fixed interaction strength, there is a critical temperature below which conductivity vanishes (the increase in temperature opens more channels and therefore even a small interaction is sufficient to delocalize the system). Their analysis of perturbation theory (in the Keldysh formalism) of course is not exact. Some diagrams have to be thrown away for the remaining ones to be tractable, and the ones that get re-summed are essentially those without loops (they are called rainbow diagrams), an approximation which is called by the authors Self-Consistent Born Approximation (SCBA, which, when further supplemented by the negligence of the real part of the self-energies, gets renamed Imaginary SCBA, or ImSCBA). This is analogous to the “upper bound" condition of [27] and has been studied, under the name of Forward Scattering Approximation, for a variety of one-electron problems on lattices and also for some many-body problems [485]. For the one-electron Anderson problem there is no doubt that this approximation provides an overestimate of the localization critical disorder, namely, the system could be localized even for smaller disorder but cannot be delocalized for larger disorder. This can be explained in a number of ways (see for example [472] for one based on the properties of the Abou-Chacra, Anderson, and Thouless integral equation [4]) and in the one-electron problem, as said before, it is beyond doubt. If this upper bound property would transfer tout-court to the many-body problem, there would be no doubt that an MBL phase would exists on the basis of Basko et al. analytic work. Unfortunately, this has not been proved yet, and the stability of MBL phase, defined in the following, remains an open question.
@@
Subsequent work in lattice systems mostly focused on infinite temperature states (the single-particle Hamiltonian is both lower and upper bounded), and, importantly, put the emphasis in the formulation of the MBL as a phase of matter that avoids thermalization [460, 451, 440]. Following these works, the non-ergodic behavior exists if for a generic local observable $A$,
$A_{\infty}\equiv\lim_{L,t\to\infty}\bar{A}(t)\neq\mathrm{Tr}\left[\hat{\rho}_{MC}\hat{A}\right],$ (6)
@@
and hence the system is said to be in an MBL phase. This is the definition of MBL that we employ in the present review. Conversely, if the thermalization occurs for a generic observable $A$, i.e., when (4) is satisfied, we say that the system in an ergodic phase. We note that our definition pertains to observables $A$ that are local in real space, i.e., the space in which the sites of the considered many-body system reside [^4] This aspect of the definition can be readily extended to include different types of observables, for instance, observables local in quasi-momentum space.. Moreover, using Eq. (6), the MBL is defined as a robust ergodicity breaking phenomenon for disordered systems and arbitrary initial states. Hence, our definition does not apply to fine-tuned settings such as integrable models [679] or many-body scars [575].
@@
The definition of the MBL phase (6) contains the double limit of infinite time, $t\to\infty$, and infinite system size, $L\to\infty$. The presence of both limits is essential to define the MBL phase with properties sharply different from the features of the ergodic phase. This leads to a question about the order of limits. Exact diagonalization studies provide access to eigenvalues and eigenvectors of many-body systems, thus naturally allowing to take the limit of infinite time, $t\to\infty$, first, as demonstrated in Eq. (2). This leads to a system size dependent average value of the considered observable $A_{\infty}(L)=\lim_{t\to\infty}\bar{A}(t)$ (the $L$ dependence is now explicitly indicated), which can be subsequently extrapolated to the large system size limit as $A_{\infty}=\lim_{L\to\infty}A_{\infty}(L)$, and compared with the microcanonical ensemble prediction $\mathrm{Tr}\left[\hat{\rho}_{MC}\hat{A}\right]$. In the context of disordered 1D many-body systems with local Hamiltonians, which are the focus of this review, we implicitly assume that the limit $L\to\infty$ in Eq. (6) is taken in the usual manner, i.e., by keeping the parameters of the Hamiltonian, such as the interaction amplitude, disorder strength, and tunneling amplitude, fixed. Alternative choices, allowing for scaling of the microscopic parameters with the system size $L$, may lead to interesting dynamical phenomena in the presence of long-range couplings, as we highlight in Sec. IV.3.3.
@@
The approach of sending $t\to\infty$ first, while practical for exact diagonalization studies, is impractical in experimental and numerical studies of time evolution of many-body systems. Let us consider, as an example, an ergodic system with diffusive dynamics, in which local observables attain their long time saturation values after the so-called Thouless time $t_{Th}=L^{2}/D$, where $D$ is the diffusion constant. To correctly assess that such a system is ergodic, one has to probe properties of local observables at times $t=aL^{2}$, where $a$ is a sufficiently large constant. In contrast, probing the local observables at smaller times $t\propto L$ may yield values of local observables different than their long time saturation values leading to an incorrect conclusion about lack of thermalization in the system. The time scale required to correctly assess whether the system is ergodic or MBL increases with the slowing down of the dynamics: for instance, for an ergodic system with subdiffusive dynamics characterized by a dynamical exponent $z>2$, one has to probe the dynamics at times scaling at least as $t\propto L^{z}$.
@@
The above examples show that in order to assess whether a system reaches the thermal equilibrium or not, one has to probe time evolution at time scales that diverge with system size $L$, and that the precise relation between $t$ and $L$ is fixed by the type of dynamics of the system. This property, coupled with a dramatic slow down of the dynamics of many-body systems with increasing the disorder strength, constitutes one of the main difficulties in assessing whether systems are in the ergodic or the MBL phase. [^5] Sending $L\to\infty$ followed by the $t\to\infty$ limit may not be the most natural approach in studies of ETH and MBL that aim to explore whether system thermalizes for generic (physically realizable) conditions. In particular, initial states of domain wall type involve lengthscales that are divergent in the $L\to\infty$ limit. This leads to divergent timescales of observables located away from the domain walls, preventing one from verifying (4) even for a diffusive system.
@@
MBL was originally defined as an insulating phase of an interacting electronic system at sufficiently low temperature [51], where conductivity and diffusion coefficients vanish. In principle, conductivity can vanish for a number of reasons, and diffusion may be substituted by sub-diffusion (in particular in one-dimensional systems), leaving the final equilibrium state to be described by the microcanonical ensemble, see also Appendix XIV. However, it was noted already in 2012 (Oganesyan, private communications) that the vanishing of conductivity in the MBL phase may be due to the existence of local integrals of motion for MBL systems (LIOMs). These were properly conjectured in [580] where also details of the implications for the MBL phenomenology was given (e.g., the slow growth of entanglement already observed in [697]), and shortly after, the perturbation series analysis of [51] [^6] Ref. [245] identified a class of terms that parametrically enhance delocalization threshold with respect to the predictions of [51, 244]. was adapted by [528] to show that indeed LIOMs should exist in the MBL phase when the same approximations of the ImSCBA are used. In parallel, numerical algorithms to find and characterize LIOMs were proposed [117, 516, 418]. In the meanwhile, a detailed mathematical analysis suggesting stability of the MBL phase of a one-dimensional disordered spin chain was published by [292], see also [293] for a parallel discussion of this work and of other perturbative approaches to MBL.
@@
Our definition of MBL phase (6) is not formulated in terms of the transport properties and it does not explicitly rely on the vanishing of the DC conductivity of the system. Those properties, however, are characteristic features of a localized system as described in the next section, and can be derived from the existence of local conserved operators, such as those that will be introduced in Eq. (9). While the perturbative treatment of [51] assumes that the given interacting system can be smoothly connected to an Anderson localized non-interacting model, our definition allows for MBL in a larger class of many-body systems, including for example systems which do not have a natural non-interacting limit. And, while the ergodicity breaking condition (6) explicitly involves the energy of the initial state, present day studies of MBL typically focus on the limit of infinite temperature analyzing eigenstates in the middle of many-body spectrum. Let us also note that the term “many-body localization” to refer to Anderson localization beyond the one-electron approximation was already in use in the 1980’s, the earliest occurrence we could found is in a paper by W. McMillan dated 1981 [407]. That paper is a generalization of the “gang of four” renormalization group analysis [5], including both conductivity and interaction in a two-parameter scaling set of RG equations (the author seems unaware of the paper [208] that was published shortly before that). As a side remark, the failure of a simple one-parameter scaling theory for describing the MBL transition was recognized in recent works and is discussed in more detail in Sec. IV.3.
@@
Nevertheless, we stress that a rigorous definition of MBL is still a matter of debate, as also highlighted in Sec. IV.1. For example, while the above discussion implicitly assumes the necessity of LIOMs in the MBL phase, this requirement may be too restrictive. Another, much stronger form of ergodicity breaking is localization in Fock space graph, also referred to as Fock space localization [164] [^7] Note that Fock space localization considered here is different from the notion of Fock space localization invoked in [51], see also Sec. VI.2.1.. In our definition here, MBL includes, but is not limited to, Fock space localization, since (6) may be valid also when many-body eigenstates are not localized in the Fock space [156, 401].

### v3

@@ III.2 Ergodicity breaking: single particle systems and beyond
While ETH has been observed numerically in a wide class of many-body systems, there is no proof of it for a generic, local Hamiltonian, and provable counterexamples of fine-tuned, integrable, Hamiltonians can be constructed. In this sense, the situation is similar to the classical mechanics’ ergodic hypothesis, which is widely believed to be true for generic Hamiltonian dynamics (for a sufficiently large number of degrees of freedom), but could only be proved for some realistic systems [31, 656].
Anderson localization [29] is often referred to as a counter-example to ETH, which has no counterpart in classical mechanics. However, this interpretation requires some care, since the Anderson model, invoked to model Anderson localization on a lattice, is a quadratic model, [^5] Quadratic Hamiltonians are sums of bilinear forms of creation and annihilation operators, see Eq. (5) for one example. and it can be proven that many-body eigenstates of quadratic models can not exhibit ETH in the many-particle Hilbert space [786]. However, recent work has formulated the notion of single-particle eigenstate thermalization [787]666 We note that observables in integrable systems relax to the values predicted by the generalized Gibbs ensemble [567] in a process called generalized thermalization [111, 734]., which can be seen as the analogue of (3), defined for the matrix elements in single-particle eigenstates of quadratic models, and is a natural consequence of the Berry conjecture [65]. Within this framework, Anderson insulator should be seen as an example of the breakdown of single-particle eigenstate thermalization.
@@
where $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) are fermionic creation (anihilation) operators creating particle at site $i$, the sum over neighboring lattice sites is denoted by $\braket{i,j}$, and $\epsilon_{i}$ are on-site energies. In case the on-site energies $\epsilon_{i}$ are random uncorrelated numbers, the model in Eq. (5) is referred to as the Anderson model. If the variance of $\epsilon_{i}$ is large enough (with respect to the hopping rate), it turns out that each particle stays confined close to its original position, and the system avoids the process of reaching thermal equilibrium. It goes without saying that there is no conduction in the system and the system is an insulator. In this case, the single-particle version of the ETH ansatz (3) does not apply [787], and the long-time value of an observable such as site occupation is not recovered by the microcanonical average (4). This gives rise to the absence of transport and several mathematical proofs of localization at large disorder exist [250, 227, 12, 88] [^7] It is curious, however, that no proof of absence of localization at small disorder exist for the Anderson model except for the Bethe lattice [13]..
In spite of the above arguments, it could not be ruled out that interactions would destroy the one-electron picture of Anderson localization and a small conductivity will be born, possibly even exponentially small in the inverse of the interaction strength (measured in units of the local density of single-particle states). In fact, a mechanism in which a small interaction between electrons and phonons can induce a non-zero conductivity via the phonon-assisted hopping was already pointed out by Mott [468]. However, crucially, the phonons considered by Mott are supposed to be a good, ergodic, bath not perturbed by the presence of localized electrons. But can a bath of localized electrons provide the necessary frequencies for a conduction probe electron to move for macroscopic distances? In other words, could electrons act like Mott’s phonon bath? Fleishman and Anderson [221] already in the 1980’s proposed, on the basis of perturbation theory (and some general intuition), that, at small enough temperatures and weak interactions, localization would be preserved.
@@ III.3 MBL phase: a definition
Twenty six years after Fleishman and Anderson, [54, 260] presented a thorough analysis of the perturbation theory, where they found that at fixed interaction strength, there is a critical temperature below which conductivity vanishes (the increase in temperature opens more channels and therefore even a small interaction is sufficient to delocalize the system). Their analysis of perturbation theory (in the Keldysh formalism) of course is not exact. Some diagrams have to be thrown away for the remaining ones to be tractable, and the ones that get re-summed are essentially those without loops (they are called rainbow diagrams), an approximation which is called by the authors Self-Consistent Born Approximation (SCBA, which, when further supplemented by the negligence of the real part of the self-energies, gets renamed Imaginary SCBA, or ImSCBA). This is analogous to the “upper bound" condition of [29] and has been studied, under the name of Forward Scattering Approximation, for a variety of one-electron problems on lattices and also for some many-body problems [524]. For the one-electron Anderson problem there is no doubt that this approximation provides an overestimate of the localization critical disorder, namely, the system could be localized even for smaller disorder but cannot be delocalized for larger disorder. This can be explained in a number of ways (see, for example, [510] for one based on the properties of the Abou-Chacra, Anderson and Thouless integral equation [4]) and in the one-electron problem, as said before, it is beyond doubt. If this upper bound property would transfer tout-court to the many-body problem, there would be no doubt that an MBL phase would exists on the basis of Basko et al. analytic work. Unfortunately, this has not been proved yet, and the stability of MBL phase, defined in the following, remains an open question.
@@
Subsequent work in lattice systems mostly focused on the infinite temperature states, and, importantly, emphasized the formulation of the MBL as a phase of matter that avoids thermalization [498, 489, 478]. Following these works, the dynamics generated by Hamiltonian $\hat{H}$ of an isolated many-body system is non-ergodic if there exists a local [^8] A local observable acts non-trivially at a finite number of lattice sites independent of the system size $L$. observable $\hat{A}$ for which
$A_{\infty}\equiv\lim_{L\to\infty}\lim_{t\to\infty}\bar{A}(t)\neq\mathrm{Tr}\left[\hat{\rho}_{MC}\hat{A}\right].$ (6)
@@
An isolated interacting many-body system is said to be in an MBL phase when the non-ergodic behavior is caused by the disorder and occurs for a robust class of initial states and observables. This is the definition of MBL that we employ in the present review. Conversely, if the thermalization occurs for all local observables $\hat{A}$, i.e., when (4) is satisfied, we say that the system in an ergodic phase. In systems in which the MBL phase exists, it is separated from the ergodic phase by an MBL phase transition. We note that our definition pertains to observables $\hat{A}$ that are local in real space, i.e., the space in which the sites of the considered many-body system reside [^9] This aspect of the definition can be readily extended to include different types of few-body observables, for instance, observables local in quasi-momentum space.. Moreover, the MBL phase is defined as a robust ergodicity breaking phenomenon for disordered systems and arbitrary physically relevant [^10] Studies of thermalization predominantly focus on product or low-entangled initial states. initial states. Hence, our definition does not apply to fine-tuned settings such as integrable models [734] or many-body scars [622].
In ergodic many-body systems, the ETH ansatz (3) is satisfied, in the large system size limit, by all eigenstates [351]. In contrast, a weak ETH allows for (i) polynomial decay with system size $L$ of the fluctuations of matrix elements, as in integrable models [315, 17, 785], and (ii) the existence of special eigenstates, dubbed outliers, which do not approach the corresponding microcanonical averages, such as quantum scars, c.f. Sec. X.7. The weak ETH is not expected to apply in the MBL phase. Instead, in the MBL phase, one anticipates that the fluctuations of matrix elements of local observables do not decay with $L$, making the ETH invalid for the vast majority of eigenstates. In that case, Eq. (6) is satisfied, in the limit $t\to\infty$ and $L\to\infty$, for the physically relevant initial states.
@@
The definition of the MBL phase (6) contains the double limit of infinite time, $t\to\infty$, and infinite system size, $L\to\infty$. The presence of both limits is essential to define the MBL phase with properties sharply different from the features of the ergodic phase. This leads to a question about the order of limits. Exact diagonalization studies provide access to eigenvalues and eigenvectors of many-body systems, thus naturally allowing to take the limit of infinite time, $t\to\infty$, first, as demonstrated in Eq. (2). This leads to a system size dependent average value of the considered observable, $A_{\infty}(L)=\lim_{t\to\infty}\bar{A}(t)$ (the $L$ dependence is now explicitly indicated), which can be subsequently extrapolated to the large system size limit as $A_{\infty}=\lim_{L\to\infty}A_{\infty}(L)$, and compared with the microcanonical ensemble prediction, $\mathrm{Tr}\left[\hat{\rho}_{MC}\hat{A}\right]$. In the context of disordered 1D many-body systems with local Hamiltonians, which are the focus of this review, we implicitly assume that the limit $L\to\infty$ in Eq. (6) is taken in the usual manner, i.e., by keeping the parameters of the Hamiltonian, such as the interaction amplitude, disorder strength and tunneling amplitude, fixed. Alternative choices, allowing for scaling of the microscopic parameters with the system size $L$, may also lead to interesting dynamical phenomena, such as, e.g., in the presence of long-range couplings, as we highlight in Sec. IV.3.3.
@@
The approach of sending $t\to\infty$ first, natural when eigenvalues and eigenvectors of $\hat{H}$ are computed, is impractical in experimental and numerical studies of time evolution of many-body systems [^11] In a solid-state physics experimental context, one usually assumes the limit $L\to\infty$ first, assuming the system to be effectively in the thermodynamic limit. One then is interested in the time evolution of local perturbation acting on a spatial region $\Omega$ of size $O(1)$ of an initially homogeneous state, as for example the creation of a particle on top of an homogeneous state or a particle-hole excitation created by a photon or a neutron scattered. If the local perturbation eventually (i.e. in the limit $t\to\infty$) escapes the region $\Omega$, then the system is not localized (neither Anderson nor MBL).. Let us consider, as an example, an ergodic system with diffusive dynamics, in which local observables attain their long time saturation values after the so-called Thouless time $t_{Th}=L^{2}/D$, where $D$ is the diffusion constant. To correctly assess that such a system is ergodic, one has to probe properties of local observables at times $t=aL^{2}$, where $a$ is a sufficiently large constant assuring that the diffusive relaxation processes are finished. In contrast, probing the local observables at smaller times $t\propto L$ may yield values of local observables different than their long time saturation values leading to an incorrect conclusion about lack of thermalization in the system. The time scale required to correctly assess whether the system is ergodic or MBL increases with the slowing down of the dynamics: for instance, for an ergodic system with subdiffusive dynamics characterized by a dynamical exponent $z>2$, one has to probe the dynamics at times scaling at least as $t\propto L^{z}$.
@@
The above examples show that in order to assess whether a system reaches the thermal equilibrium or not, one has to probe time evolution at times scales that diverge with system size $L$, and that the precise relation between $t$ and $L$ is fixed by the type of dynamics of the system. This property, coupled with a dramatic slow down of the dynamics of many-body systems with increasing the disorder strength, constitutes one of the main difficulties in assessing whether systems are in the ergodic or the MBL phase. [^12] Sending $L\to\infty$ followed by the $t\to\infty$ limit may not be the most natural approach in studies of ETH and MBL that aim to explore whether system thermalizes for generic (physically realizable) conditions. In particular, initial states of domain wall type involve lengthscales that are divergent in the $L\to\infty$ limit. This leads to divergent timescales of observables located away from the domain walls, preventing one from verifying (4) even for a diffusive system.
@@
MBL was originally defined as an insulating phase of an interacting electronic system at sufficiently low temperature [54], where conductivity and diffusion coefficients vanish. In principle, conductivity can vanish for a number of reasons, and diffusion may be substituted by sub-diffusion (in particular in one-dimensional systems), leaving the final equilibrium state to be described by the microcanonical ensemble, see also Appendix XIV. However, it was noted already in 2012 (Oganesyan, private communications) that the vanishing of conductivity in the MBL phase may be due to the existence of local integrals of motion for MBL systems (LIOMs). These were properly conjectured in [627] where also details of the implications for the MBL phenomenology was given (e.g., the slow growth of entanglement already observed in [753]), and shortly after, the perturbation series analysis of [54] [^13] Ref. [261] identified a class of terms that parametrically enhance delocalization threshold with respect to the predictions of [54, 260]. was adapted by [574] to show that, indeed, LIOMs should exist in the MBL phase when the same approximations of the ImSCBA are used. In parallel, numerical algorithms to find and characterize LIOMs were proposed [122, 557, 454]. In the meanwhile, a detailed mathematical analysis suggesting stability of the MBL phase of a one-dimensional disordered spin chain was published by [316, 317], see also [318] for a parallel discussion of this work and of other perturbative approaches to MBL. Starting from the analysis of [317], refined, and purged from a hypothesis about absence of level attraction, the authors of the recent paper [573] claim a proof of the existence of LIOMs in a fraction of samples which is exponentially small in system size. Although this is not yet a proof of the existence of an MBL phase, using this result, the authors can prove the absence of diffusion in a strongly disordered interacting quantum spin-1/2 chain, namely the disordered quantum Ising model introduced in Eq. (87), leading to the vanishing of DC conductivity and of any diffusion coefficient.
@@
Our definition of the MBL phase (6) is not formulated in terms of the transport properties and it does not explicitly rely on the vanishing of the DC conductivity of the system. Those properties, however, are characteristic features of a localized system as described in the next Section, and can be derived from the existence of local conserved operators, such as those that will be introduced in Eq. (9). While the perturbative treatment of [54] assumes that the given interacting system can be smoothly connected to an Anderson localized non-interacting model, our definition allows for MBL in a larger class of many-body systems, including for example systems which do not have a natural non-interacting limit. And, while the ergodicity breaking condition (6) explicitly involves the energy of the initial state, present day studies of MBL typically focus on the limit of infinite temperature analyzing eigenstates in the middle of many-body spectrum. Let us also note that the term “many-body localization” to refer to Anderson localization beyond the one-electron approximation was already in use in the 1980’s, the earliest occurrence we could found is in a paper by W. McMillan dated 1981 [443]. That paper is a generalization of the “gang of four” renormalization group (RG) analysis [5], including both conductivity and interaction in a two-parameter scaling set of RG equations (the author seems unaware of the paper [221] that was published shortly before that). As a side remark, the failure of a simple one-parameter scaling theory for describing the MBL transition was recognized in recent works and is discussed in more detail in Sec. IV.3.
@@
Nevertheless, we stress that a precise definition of MBL is still a matter of debate, as also highlighted in Sec. IV.1. For example, while the above discussion implicitly assumes the necessity of LIOMs in the MBL phase, this requirement may be too restrictive. Another, stronger form of ergodicity breaking is localization in Fock space graph, also referred to as Fock space localization [171] [^14] Note that Fock space localization considered here is not entirely identical to the notion of Fock space localization invoked in [54], see also Sec. VI.2.1.. In our definition here, MBL includes, but is not limited to, Fock space localization, since (6) may be valid also when many-body eigenstates are not localized in the Fock space [163, 437].